OpenStudy (anonymous):

Solve 2x^2 + 20x = –38. Show your work please.

jhonyy9 (jhonyy9):

x_1,2=(-20+/- sqrt(400-304))/4 =(-20+/- sqrt96)/4 =(-20+/- sqrt(16*6))/4 = =(-20+/- 4sqrt6)/4 = -5+/- sqrt6